OpenStudy (anonymous):

The radio or the earth is approximately of 6 370 Km. A 20-Kg object is rised to a 160Km over the earth surface. a) what is the weight of that object at that height?

OpenStudy (anonymous):

g' =g(1 - 2h/R)

OpenStudy (anonymous):

w=mg'

OpenStudy (anonymous):

Thanks, now im working with the numbers in oder to get the answer ...

OpenStudy (anonymous):

should I tranform the units? coz i dont get the answer in my book. g = 9.8; h = 160; R = 797; m = 20; g1 = g (1 - 2 h/R); W = m*g1 W=117.305 the answer in my book is 186.5N.

OpenStudy (anonymous):

then i think i got it wrng.. g` = g(1 - h/R) srry

OpenStudy (anonymous):

No. I was wrong replacing the numbers lol